Accountancy

Rama, brothers earn a net profit of Rs 25000 with a capital of Rs 200000.The normal rate of return in the business is 10% . Use capitalization of super profit method to value the goodwill
Answer #1

Normal profit=200000×10/100

Super profit=average profit- normal profit 

=25000-20000=5000

Goodwill=super profit×100/normal rate of return 

5000×100/10=50000

answered by: Advika